# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

### v1

@@ Abstract
Turbulent flows, despite their apparent randomness, exhibit coherent structures that underpin their dynamics. Proper orthogonal decomposition (POD) has been widely used to extract these structures from experimental data. While periodic features like vortex shedding can be identified using POD mode pairs when periodicity dominates the flow, detecting such structures in complex flows is more challenging. The Hilbert proper orthogonal decomposition (HPOD) addresses this by applying POD to the analytic signal of the turbulent fluctuations, yielding complex modes with a $90^{\circ}$ phase shift between the real and imaginary components. These modes capture propagating structures effectively but introduce filtering artefacts from the Hilbert transform that is used to derive the analytic signal. The current work investigates the relationship between the modes of the POD and those of the HPOD on the velocity fluctuations in the wake of a sphere. By comparing their outputs, POD mode pairs that correspond to the same propagating structures revealed by HPOD are identified. Furthermore, this study explored whether computing the analytic signal of the POD modes can replicate the HPOD modes, offering a more computationally efficient method for determining the pairs of POD modes that represent propagating structures. The results show that the pairs of POD modes identified by the HPOD can be more efficiently determined using the Hilbert transform directly on the POD modes. This method enhances the interpretive power of POD, enabling more detailed analysis of turbulent dynamics without introducing the filtering from the Hilbert transform.
@@ 1 Introduction
While turbulence often appears random, many turbulent flows contain underlying structures, such as eddies and vortices, the scale and nature of which vary depending on the conditions of the flow. Identifying these structures allows for a more detailed description of the flow, as well as the isolation of the dominant structures in the turbulence based on their contribution to the quantity or phenomenon of interest, such as turbulent kinetic energy (TKE) or mixing. Ranking these structures based on their significance to a given application or study allows the construction of parsimonious reduced-order models, which can be used to reduce the computational intensity of analyses or to reduce noise from experimental data.
@@
The complex flow phenomena present in the flow over bluff bodies, even at low Reynolds numbers, develop as the Reynolds number increases, and are a subject of ongoing study. In this respect, the flow over spheres serves as a canonical model for flows over axisymmetric bluff bodies, and is of particular interest for developing a broad understanding of these flow phenomena. The phenomena observed in the wake of a sphere include transition from axisymmetry to planar symmetry to asymmetry, the formation and shedding of vortices, the development of shear layer instabilities, and flow separation. As the flow over a sphere varies significantly with Reynolds number, experimental and numerical investigations have covered a wide range of Reynolds number regimes.
@@
In the context of experimental studies of the flow over spheres, one of the main challenges is fixing the sphere in a wind or water tunnel test section with minimal disturbance of the flow, since more often than not, the mounting/support structure of the sphere can have a significant influence on the flow. The supports chosen for an experiment should simultaneously provide rigid and stable positioning of the sphere, and minimise the disturbance to the flow, while also facilitating the placement of probes without obstructing optical measurement access. Some of the supports used in these studies include rigid downstream supports [1], honeycomb structures with rigid supports upstream of the sphere [2, 3], and wires placed azimuthally [4, 5]. The drag on the spheres has been measured using hot-wire probes [1, 6], with flow visualisations performed using aluminium dust [7], smoke [8, 5], and dyes [9, 4, 10] to gain a qualitative understanding of the recirculation region, flow separation, and vortex shedding in the wake. Various forms of particle image velocimetry (PIV), including planar [5], stereographic [3] and tomographic [11], have been used to quantitatively measure the flow around spheres.
@@
While mounting the sphere is not a concern in numerical studies, simulations are subject to their own set of challenges. The geometry of the problem and the range of scales involved require careful consideration of the mesh and solver to resolve the flow at high Reynolds numbers. Because of this complexity, large eddy simulation (LES[12, 13, 14, 15] and detached eddy simulation (DES[14, 16] methods have been favoured over the more computationally expensive direct numerical simulation (DNS[12, 15, 17, 18] to resolve the main features of the flow at high Reynolds numbers. In conjunction with the advancement of computational capabilities, these models have extended the feasibility of numerical investigations of the flow over spheres to large Reynolds numbers.
@@
Experimental and numerical studies have examined the development of the flow over spheres over a wide range Reynolds number regimes. At low Reynolds number, $Re_{D}<20$, the flow in the wake is laminar and axisymmetric, and there is no recirculation [7, 13]. A small vortex ring forms near the rear stagnation point around $Re_{D}\approx 25$ [7], and elongates in the streamwise direction with increasing Reynolds number. This vortex ring begins to oscillate around $Re_{D}\approx 130$ with a long period, which becomes more prominent at higher Reynolds numbers. The recirculation region formed behind the sphere, which exhibits reverse flow in the mean, extends further downstream of the sphere as Reynolds number increases [9].
@@
The axisymmetry in the wake at low Reynolds number is lost due to the formation of a double-threaded wake around $Re_{D}\approx 210$ [9, 13], which consists of two streamwise vortices of opposing sign that can be seen as dye threads in the wake of a free-falling sphere [19]. Single-frequency vortex shedding begins around $Re_{D}\approx 300$ [7, 4], and hairpin-shaped vortices form at $Re_{D}\approx 400$ [10]. The generation of an unsteady random side force, which has been observed in both experimental [8] and numerical [20] studies results in the loss of planar symmetry in the wake in the range of $400<Re_{D}<500$ [7].
@@
From $Re_{D}=800$, the small-scale Kelvin-Helmholtz instability in the shear layer at the edge of the recirculation region manifests as axisymmetric vortex shedding [16, 21]. This is the higher frequency of two instability modes identified in the wake of the sphere by [22], with the lower frequency mode arising from large-scale vortex shedding. Small-scale vortical structures are shed from the main vortical structure around $Re_{D}\approx 1000$ [20], and the recirculation region behind the sphere begins to shrink with increasing Reynolds number. Periodic fluctuations are present in the wake for much higher Reynolds numbers, $Re_{D}=O(10^{6})$ [10].
@@
Since its first application to turbulent flows by [23], proper orthogonal decomposition (POD) has been widely used for identifying and analysing structures in turbulent flows [24, 25]. As POD does not require temporal resolution of the flow and can thus be applied to non-time-resolved data, it is a robust and simple method for identifying the structures within a turbulent flow. In the case of the flow over a cylinder, the turbulent fluctuations due to the vortices shed from the sphere can be identified in the leading POD modes [26, 27]. The energy of these modes is significantly larger than the higher-order modes, and their phase angle is clearly related to that of the pressure fluctuations resulting from the shed vortices [28]. Similarly, periodic structures in screeching, impinging and free jets [29, 30, 31], are represented by the leading POD modes. Identifying pairs of POD modes which represent periodic structures allows for the a posteriori phase-averaging of experimental data which is collected asynchronously or at low frequency relative to these structures. However, the complexity of the flow over a sphere in this regime makes identification of periodic structures in the turbulent fluctuations challenging, as POD modes representing different phases of a structure are not linked by the decomposition.
@@
The analytic signal of the turbulent fluctuations, which is defined using the Hilbert transform, can be decomposed into complex modes, which include phase information, using POD [32]. While the Hilbert transform is performed in time for time-resolved data, it can also performed in an analogous spatial direction when the data is not time-resolved [33, 34, 35]. This results in modes emphasising structures which are periodic in the chosen direction, with the streamwise direction being the most sensible choice to identify structures which propagate downstream from the sphere. Although a useful method for identifying propagating structures, Hilbert POD (HPOD) introduces artefacts to the modes as a result of the Hilbert transform.
@@
The current work applies both POD and HPOD, with the Hilbert transform performed in the streamwise direction, to experimentally-measured velocity fluctuations in the wake of a sphere at $Re_{D}=7780$. The modes resulting from each of these decompositions are compared, and the correlation between the POD modes and the HPOD modes is used to determine pairs of POD modes representing propagating structures. The data is then phase-averaged using the elucidated POD mode pairs to observe the periodic behaviour of the propagating structures. By pairing POD modes using the HPOD modes, the dynamics of the turbulent structures within the wake can be determined while avoiding the filtering that is introduced by the Hilbert transform. Additionally, the same mode pairing can be achieved using the streamwise Hilbert transform of the individual POD modes. This method significantly reduces the computational cost of identifying paired POD modes, as it does not require POD to be performed on the complex-valued analytic signal. It is shown that the leading modes of POD and HPOD are similar in shape, and have comparable phase angles throughout the experimental data. However, the HPOD modes accentuate the propagating parts of the structures and introduce artefacts which are not present in the POD modes.
@@ 2.1 Proper Orthogonal Decomposition
POD decomposes statistically independent realisations of the random field $\boldsymbol{f}(\boldsymbol{x},t)$ into orthonormal spatial modes $\psi_{k}(\boldsymbol{x})$ with temporal coefficients $a_{k}(t)$, which are ranked by their energy $\lambda_{k}$, from largest to smallest, such that
@@
and is analogous to singular value decomposition (SVD) and principle component analysis (PCA). In the application of POD to turbulence, $\boldsymbol{f}(\boldsymbol{x},t)$ represents the turbulent fluctuations, and $\lambda_{k}$ represents the contribution of the $k^{th}$ mode to the total TKE of $\boldsymbol{f}(\boldsymbol{x},t)$. A common use for POD is the definition of a reduced-order representation of $\boldsymbol{f}(\boldsymbol{x},t)$ by reconstructing the original data using $\hat{K}<K$ modes
@@
[…]
is calculated, where $T$ is the temporal domain. The spatial modes and the corresponding energy are given by the eigendecomposition of $C(\boldsymbol{x},\boldsymbol{x}^{\prime})$
[…]
where $S$ is the spatial domain. The temporal coefficients are the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $\psi(\boldsymbol{x})$
[…]
Alternatively, POD can be performed using the method of snapshots, which first determines the temporal coefficients, and then the spatial modes. This method is more computationally efficient when the number of spatial coordinates included in $\boldsymbol{f}(\boldsymbol{x},t)$ is larger than the number of realisations of the field. The two-point correlation in time
[…]
is used instead of that in space, given in equation (4). The eigendecomposition of $C(t,t^{\prime})$
[…]
yields the eigenvalues and temporal coefficients. The spatial modes are then determined by the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $a(t)$
[…]
The temporal coefficients of the $i^{th}$ and $j^{th}$ POD modes can be expressed in terms of their phase angle $\phi_{i,j}$ as
$a_{i}(t)=\sqrt{2\lambda_{i}}\cos{\big{(}\phi_{i,j}(t)\big{)}},$ (10)
@@
$a_{j}(t)=\sqrt{2\lambda_{j}}\sin{\big{(}\phi_{i,j}(t)\big{)}},$ (11)
respectively. Thus, the phase angle $\phi_{i,j}(t)$ is given by
@@
The phase average of the data using the $i^{th}$ and $j^{th}$ POD modes at the phase angle $\phi$ is given by
@@
where $\phi_{i,j}(t)$ is defined by equation (12). While any two POD modes can be used for phase-averaging $\boldsymbol{f}(\boldsymbol{x},t)$, not all selections of modes result in physically meaningful phase averages. In the case of strongly dominant periodic fluctuations, the temporal coefficients of the corresponding modes form a circle [26]. Large periodic structures can also be identified by comparing the phase angle of the modes to an independent measurement of the phase of the flow, such as pressure signals [28]. However, the less prominent the periodic structures, the more difficult it is to relate the POD modes which represent them.

### v2

@@ Abstract
Turbulent flows, despite their apparent randomness, exhibit coherent structures that underpin their dynamics. Proper orthogonal decomposition (POD) has been widely used to extract these structures from experimental data. Periodic features such as vortex shedding can appear as POD mode pairs in strongly periodic flows, but detecting propagating structures in more complex flows is challenging. Hilbert proper orthogonal decomposition (HPOD) addresses this by applying POD to the analytic signal of the turbulent fluctuations, which yields complex modes with a $\pi/2$ phase shift between the real and imaginary components. These modes capture propagating structures effectively but introduce spectral leakage from the Hilbert transform used to derive the analytic signal. The current work investigates the relationship between the modes of the POD and those of the HPOD on the velocity fluctuations in the wake of a sphere. By comparing their outputs, POD mode pairs that correspond to the same propagating structures revealed by HPOD are identified. Furthermore, this study explores whether computing the analytic signal of the POD modes can replicate the HPOD modes, offering a more computationally efficient method for determining the pairs of POD modes that represent propagating structures. The results show that the pairs of POD modes identified by the HPOD can be determined more efficiently using the Hilbert transform directly on the POD modes. This method enhances the interpretive power of POD, enabling more detailed analysis of turbulent dynamics without the need to compute the analytic signal of the entire turbulent fluctuation data.
Keywords: Wakes, Turbulent flows, Low-dimensional models
@@ 1 Introduction
While turbulence often appears random, many turbulent flows contain underlying structures, such as eddies and vortices. The scale and nature of these structures vary depending on the flow conditions. Identifying these structures allows for a more detailed description of the flow, as well as the isolation of the dominant structures within the turbulence based on their contribution to the quantity or phenomenon of interest, such as turbulent kinetic energy (TKE) or mixing. Ranking these structures based on their significance to a given application or study allows the construction of parsimonious reduced-order models, which can be used to reduce the computational intensity of analyses or to reduce noise from experimental data.
@@
The complex flow phenomena present in the flow over bluff bodies, even at low Reynolds numbers, develop with increasing Reynolds number, and remain a subject of ongoing study. In this respect, the flow over spheres serves as a canonical model for flows over axisymmetric bluff bodies, and is of particular interest for advancing the understanding of these flow phenomena. The phenomena observed in the wake of a sphere include transition from axisymmetry to planar symmetry, from planar symmetry to asymmetry, the formation and shedding of vortices, the development of shear layer instabilities, and flow separation. As the flow over a sphere varies significantly with Reynolds number, experimental and numerical investigations have explored a wide range of Reynolds number regimes.
@@
In the context of experimental studies of the flow over spheres, one of the main challenges is fixing the sphere in a wind or water tunnel test section with minimal disturbance of the flow, since the mounting or support structure of the sphere frequently has a significant influence on the flow. The supports chosen for an experiment should simultaneously provide rigid and stable positioning of the sphere, and minimise the disturbance to the flow, while facilitating probe placement and maintaining unobstructed optical access for measurements. Some of the supports used in these studies include rigid downstream supports [1], honeycomb structures with rigid supports upstream of the sphere [15, 23], and wires arranged azimuthally [34, 17]. The drag on the spheres has been measured using hot-wire probes [1, 44], with flow visualisations performed using aluminium dust [38], smoke [39, 17], and dyes [18, 34, 2] to provide a qualitative understanding of the recirculation region, flow separation, and vortex shedding in the wake. Various forms of particle image velocimetry (PIV), including planar [17], stereoscopic [23], and tomographic [12], have been used to quantitatively measure the flow around spheres.
@@
While mounting the sphere is not a concern in numerical studies, simulations are subject to distinct challenges. The problem geometry and the wide range of scales involved require careful consideration of the mesh and solver to accurately resolve the flow at high Reynolds numbers. Because of this complexity, large eddy simulation (LES[41, 42, 9, 33] and detached eddy simulation (DES[9, 10] methods have been preferred over the more computationally expensive direct numerical simulation (DNS[41, 33, 31, 32] for resolving the primary features of the flow at high Reynolds numbers. In conjunction with the advancement of computational capabilities, these models have expanded the scope of numerical investigations of the flow over spheres to large Reynolds numbers.
@@
Experimental and numerical studies have examined the development of the flow over spheres over a wide range of Reynolds number regimes. At low Reynolds number, $Re_{D}<20$, the flow in the wake is laminar and axisymmetric, and recirculation is absent [38, 42]. A small vortex ring forms near the rear stagnation point around $Re_{D}\approx 25$ [38], and elongates in the streamwise direction with increasing Reynolds number. This vortex ring begins to oscillate around $Re_{D}\approx 130$ with a long period, which becomes more prominent at higher Reynolds numbers. The recirculation region formed behind the sphere, which exhibits reverse flow in the mean flow, extends further downstream of the sphere as Reynolds number increases [18].
@@
The axisymmetry in the wake at low Reynolds number is lost due to the formation of a double-threaded wake around $Re_{D}\approx 210$ [18, 42], which consists of two streamwise vortices of opposite sign that can be seen as dye threads in the wake of a free-falling sphere [22]. Single-frequency vortex shedding begins around $Re_{D}\approx 300$ [38, 34], and hairpin-shaped vortices form at $Re_{D}\approx 400$ [2]. The generation of an unsteady random side force, which has been observed in both experimental [39] and numerical [49] studies, leads to the loss of planar symmetry in the wake within the range of $400<Re_{D}<500$ [38].
@@
For $Re_{D}\gtrapprox 800$, the small-scale Kelvin-Helmholtz instability in the shear layer at the edge of the recirculation region appears as axisymmetric vortex shedding [10, 8]. This is the higher-frequency of two instability modes identified in the wake of the sphere by Kim & Durbin [19], with the lower-frequency mode arising from large-scale vortex shedding. Small-scale vortical structures are shed from the main vortical structure around $Re_{D}\approx 1000$ [49], and the recirculation region behind the sphere begins to shrink with increasing Reynolds number. Periodic fluctuations are present in the wake for significantly higher Reynolds numbers, $Re_{D}=O(10^{6})$ [2].
@@
Since its first application to turbulent flows by Lumley [21], proper orthogonal decomposition (POD) has been widely used for identifying and analysing structures in turbulence [35, 4]. As POD does not require temporal resolution of the flow and can thus be applied to non-time-resolved data, it is a robust yet straightforward method for identifying the structures within a turbulent flow. In the case of the flow over a cylinder, the turbulent fluctuations due to the vortices shed from the cylinder can be identified in the leading POD modes [24, 26]. The energy of these modes is significantly larger than the higher-order modes, and their phase angle is clearly related to that of the pressure fluctuations resulting from the shed vortices [25]. Similarly, periodic structures in screeching, impinging, and free jets [13, 45, 46] are represented by the leading POD modes. Identifying pairs of POD modes which represent periodic structures enables the a posteriori phase-averaging of experimental data collected asynchronously or at a low sampling frequency relative to these structures. However, the complexity of the flow over a sphere in this regime makes identification of periodic structures in the turbulent fluctuations challenging, as POD modes representing different phases of a structure are not inherently linked by the decomposition.
@@
The analytic signal of the turbulent fluctuations, which is defined using the Hilbert transform, can be decomposed into complex modes that include phase information using POD [16]. While the Hilbert transform is performed in time for time-resolved data, it can also be performed in an analogous spatial direction when the data is not time-resolved [27, 28, 20]. This results in modes emphasising structures which are periodic in the chosen direction, with the streamwise direction being the most appropriate choice to identify structures which propagate downstream from the sphere. Although a useful method for identifying propagating structures, Hilbert POD (HPOD) introduces spectral leakage to the modes due to the Hilbert transform being derived from the Fourier transform, which inherently assumes periodicity.
Raiola & Kriegseis [29] demonstrated the ability of spatial HPOD to extract spatiotemporally coherent wavepackets from flows with significant advecting structures. Their application of HPOD to DNS of a laminar vortex street in the wake of a cylinder at $Re_{D}=100$ showed that the spatial HPOD provides a more compact decomposition than POD. In this case, the first two POD modes representing the vortex street were captured by a single HPOD mode. In this example, the leading mode of the spatial HPOD was able to capture the first two POD modes even when the data were shuffled in time and the dataset reduced in size, whereas the modes of the conventional time-based HPOD converged towards the POD modes, and no longer captured paired POD modes in a single HPOD mode. They further applied HPOD to LES of a turbulent jet at $Re_{D}=10^{6}$ [43], showing that the spatial HPOD could still identify advecting wavepackets in a flow with a broader range of scales. Finally, HPOD was applied to PIV data from a turbulent subsonic jet [30] at $Re=33,000$, which lacked the temporal resolution afforded by the LES data. Despite this limitation, the spatial HPOD was able to extract spatial modes similar to those identified in the LES data.
@@
The current work applies both POD and HPOD, with the Hilbert transform performed in the streamwise direction following the methodology of [29], to experimentally measured velocity fluctuations in the wake of a sphere at $Re_{D}=7780$. Modes resulting from each decomposition are compared, and correlations between the POD and HPOD modes are used to identify POD mode pairs representing propagating structures. A novel approach is introduced in which the Hilbert transform is applied directly to the POD modes, allowing mode pairs associated with periodic or propagating structures to be identified without performing POD on the complex analytic signal of the turbulent fluctuations. This method significantly reduces the computational cost of identifying propagating structures and preserves the energy ranking of the modes relative to the original data. The identified POD mode pairs are then used to phase-average the velocity fluctuations in the wake of the sphere to examine the propagating structures, while phase-averaging based on HPOD modes is also performed to demonstrate the influence of HPOD on the identified structures due to its emphasis on propagating structures and inherent assumption of periodicity.
Section 2 outlines the theoretical aspects of POD and HPOD and the rationale for applying the Hilbert transform directly to the POD modes. Section 3 provides details of the experimental facility and measurement methodology. Section 4 presents the analysis and results of the POD and HPOD, the pairing of POD modes using HPOD modes, the application of the Hilbert transform directly to the POD modes, and the resulting phase averages of the turbulent fluctuations. Section 5 summarises the work, highlights the key findings, and provides closing remarks.
@@ 2.1 Proper Orthogonal Decomposition
POD decomposes statistically independent snapshots of the random field $\boldsymbol{f}(\boldsymbol{x},t)$ into orthonormal spatial modes $\psi_{k}(\boldsymbol{x})$ with temporal coefficients $a_{k}(t)$, which are ranked by their associated energy $\lambda_{k}$ in descending order, such that
@@
and is mathematically analogous to singular value decomposition (SVD) and principal component analysis (PCA). In the application of POD to turbulence, $\boldsymbol{f}(\boldsymbol{x},t)$ represents the turbulent fluctuations, and $\lambda_{k}$ represents the contribution of the $k^{th}$ mode to the total TKE of $\boldsymbol{f}(\boldsymbol{x},t)$. A common use for POD is the definition of a reduced-order representation of $\boldsymbol{f}(\boldsymbol{x},t)$ by reconstructing the original data using $\hat{K}<K$ modes
@@
$C(\boldsymbol{x},\boldsymbol{x}^{\prime})=\int_{T}\boldsymbol{f}(\boldsymbol{x},t)\boldsymbol{f}(\boldsymbol{x}^{\prime},t)\,dt,$ (4)
is calculated, where $T$ denotes the temporal domain. The spatial modes and the corresponding energy are given by the eigendecomposition of $C(\boldsymbol{x},\boldsymbol{x}^{\prime})$
$\int_{S}C(\boldsymbol{x},\boldsymbol{x}^{\prime})\psi(\boldsymbol{x}^{\prime})\,d\boldsymbol{x}^{\prime}=\lambda\psi(\boldsymbol{x}),$ (5)
where $S$ denotes the spatial domain. The temporal coefficients are computed by projecting $\boldsymbol{f}(\boldsymbol{x},t)$ onto $\psi(\boldsymbol{x})$
$a(t)=\int_{S}\boldsymbol{f}(\boldsymbol{x},t)\psi(\boldsymbol{x})\,d\boldsymbol{x}.$ (6)
Alternatively, POD can be implemented using the method of snapshots, which first computes the temporal coefficients, and subsequently determines the spatial modes. This approach is more computationally efficient when the number of spatial coordinates in $\boldsymbol{f}(\boldsymbol{x},t)$ exceeds the number of realisations of the field. The two-instant temporal correlation
$C(t,t^{\prime})=\int_{S}\boldsymbol{f}(\boldsymbol{x},t)\boldsymbol{f}(\boldsymbol{x},t^{\prime})\,d\boldsymbol{x},$ (7)
is used instead of the spatial correlation defined in equation (4). The eigendecomposition of $C(t,t^{\prime})$
$\int_{T}C(t,t^{\prime})a(t^{\prime})\,dt^{\prime}=\lambda a(t),$ (8)
yields the eigenvalues and corresponding temporal coefficients. The spatial modes are subsequently obtained by projecting $\boldsymbol{f}(\boldsymbol{x},t)$ onto $a(t)$
$\psi(\boldsymbol{x})=\int_{T}\boldsymbol{f}(\boldsymbol{x},t)a(t)\,dt.$ (9)
When the $i^{th}$ and $j^{th}$ POD modes represent the same periodic structure separated in phase by $\pm\pi/2$, their temporal coefficients can be expressed as functions of the phase angle $\phi_{i,j}$ of the structure as
$a_{i}(t)=\sqrt{2\lambda_{i}}\cos(\phi_{i,j}(t)),$ (10)
@@
$a_{j}(t)=\sqrt{2\lambda_{j}}\sin(\phi_{i,j}(t)),$ (11)
respectively. Consequently, the phase angle $\phi_{i,j}(t)$ is given by
@@
The phase-averaged representation of the data using the $i^{th}$ and $j^{th}$ POD modes at the phase angle $\phi$ is defined as
@@
where $\phi_{i,j}(t)$ is defined by equation (12). Although any two POD modes can be employed for phase-averaging $\boldsymbol{f}(\boldsymbol{x},t)$, not all mode selections yield physically meaningful phase averages. For strongly dominant periodic fluctuations, the temporal coefficients of the corresponding modes exhibit a circular distribution in the coefficient space [24]. Large-scale periodic structures can also be identified by comparing the phase angle of the modes to an independent phase measurement of the flow, such as pressure signals [25]. However, as periodic structures become less prominent, establishing a clear relationship between the corresponding POD modes becomes increasingly challenging.